OpenStudy (anonymous):

If I varies inversely with R and I = 0.15 when R = 50, what equation should be used to show this relationship?

OpenStudy (anonymous):

A. l = 7.5/R B. l = 7.5R C. lR = 0.75 D. l = R/7.5